A sample of gas at 26°C has a pressure of 1.5 atm. Calculate the final pressure in atmospheres if the temperature is changed to 98°C at constant volume.


1
Expert's answer
2021-11-15T10:55:32-0500

Solution.

"\frac{p1}{T1} = \frac{p2}{T2}"

"p2 = \frac{p1 \times T2}{T1}"

p2 = 1.86 atm

Answer:

p2 = 1.86 atm